OpenStudy (xxjosieslayzxx):

Find the percent of the number. 522% of 85 A. 4.437 B. 44.37 C. 443.7 D. 44,370

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (ltrout):

I gave you the equation on the last question. Still remember it?

OpenStudy (ltrout):

I'll let @skullpatrol take this :)

OpenStudy (skullpatrol):

Hint: % = 1/100

OpenStudy (skullpatrol):

522% = 522/100 = 5.22

OpenStudy (skullpatrol):

In general: $$\Huge x \text{%} =\dfrac{a}{b}$$ $$\Huge 522\text{%} = \dfrac{a}{85}$$ $$\Huge 5.22•85 = a$$

OpenStudy (ltrout):

He's offline btw @skullpatrol

OpenStudy (skullpatrol):

Thanks @ltrout :-)

OpenStudy (skullpatrol):

In percent problems the word "of" means "multiply." Thus, 522% of 85 means 522/100 multiplied by 85.